OpenStudy (anonymous):

The sum of two numbers is 47 and the difference is 3 what are the numbers ? Larger Number and Smaller number

OpenStudy (anonymous):

@triciaal

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

you can do this with no algebra at all think of two numbers that add to 47 and are three apart

OpenStudy (misty1212):

for example \(20+27\) add to 47, but they are not 3 apart, lets try again

OpenStudy (misty1212):

\[25+22=47\] and they are also three apart, so \(25\) and \(22\) will do the trick

OpenStudy (misty1212):

you can also write \[x+y=47\\ x-y=3\] if you want to do it the math teacher way

OpenStudy (misty1212):

add them up in a column, get \[2x=50\] telling you that \(x=25\) so \(y=22\) but this is largely a waste of time, doing the algebra way